OpenStudy (anonymous):

x^2/x^2-4=x/x+2+2x/2-x

OpenStudy (espex):

Assuming this is your question\[\frac{x^{2}}{x^{2}-4}=\frac{x}{x+2}+\frac{2x}{2-x}\]what are you being asked to do, simplify or solve?